A grating has 10,000 lines per cm. Calculate the maximum number of principal maxima that can be formed for light of wavelength 487 nm.
1
Expert's answer
2020-02-07T10:35:51-0500

Solution. The distance between slits is 


"d=\frac {1cm}{10000}=10^{-4}cm=10^{-6}m"

Once a value for the diffraction grating’s slit spacing d has been determined, the angles for the sharp lines can be found using the equation


"d sin\theta =m\lambda"

where Θ is angle between normal and maximum direction; m is number of principal maxima; λ=487nm is wavelength. Therefore


"m=\frac{d sin\theta}{\lambda}"

Using the properties of sine function


"sin\theta \eqslantless 1"

Hense


"m\eqslantless\frac{d }{\lambda}"

"\frac{d }{\lambda} =\frac{10^{-6}}{487\times 10^{-9}}=2.053"

As result get the maximum number of principal maxima is equal to 2.


Answer. 2.